# Rigidity results for initial data sets satisfying the dominant energy condition

### v3

@@ Appendix A A metric characterization of the ball
  • The second fundamental form of $\partial M$ is positive semi-definite.
  • There exists a point in $\partial M$, where the second fundamental form is positive definite.
@@
For small enough $\varepsilon>0$, we consider the smooth manifold
@@
where $\sim$ identifies $\partial M\times(-\varepsilon,0]\subset\partial M\times(-\varepsilon,\infty)$ with an open neighborhood of $\partial M\subset M$ by the normal exponential map along $\partial M$ with respect to $g$. Denote the metric on $\partial M$ induced by $g$ by $g_{0}$ and the second fundamental form of $\partial M$ (considered as a $(1,1)$-tensor field) by $h$. The Gauss lemma, the Riccati equation for the second fundamental form and the vanishing of the curvature of $M$ imply that on $\partial M\times(-\varepsilon,0]$ the metric $g$ takes the form $g=dt^{2}+g_{t}$ where $g_{t}$ is the metric on $\partial M$ given by $g_{t}(X,Y)=g_{0}((\textup{id}+th)^{2}X,Y)$. Here $t\in(-\varepsilon,\infty)$ is the standard parameter. Since $h$ is positive semi-definite, we can use this formula to extend $g$ to $\partial M\times[0,\infty)$ and hence to a smooth metric $g^{\prime}$ on all of $M^{\prime}$. By Theorem 7.2 in [BGM05], the manifold $(M^{\prime},g^{\prime})$ is flat. Since the projection $(\partial M\times[0,\infty),dt^{2}+g_{t})\to(\partial M,g_{0})$ is $1$-Lipschitz, it is easy to see that $(M^{\prime},g^{\prime})$ is complete.
Let $\tilde{M}^{\prime}$ be the universal covering of $M^{\prime}$ equipped with the complete flat Riemannian metric $\tilde{g}^{\prime}$ induced by $g^{\prime}$. Since $(\tilde{M}^{\prime},\tilde{g}^{\prime})$ is complete, flat and simply connected, it is isometric to Euclidean space $\mathbb{R}^{n}$. Using this isometry, we regard the preimage $\tilde{M}\subset\tilde{M}^{\prime}$ of $M\subset M^{\prime}$ as a closed subset of $\mathbb{R}^{n}$ with smooth boundary $\partial\tilde{M}$. Since the embedding $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, $\tilde{M}$ is connected and $\tilde{M}$ is the universal covering of $M$.
We first assume that $n\geq 3$. By assumption, the second fundamental form of $\partial\tilde{M}\subset\mathbb{R}^{n}$ is positive semi-definite and there exists a connected component $\Sigma\subset\partial\tilde{M}$ with the property that the second fundamental form is positive definite at some point on $\Sigma$. This implies that the sectional curvature of the induced Riemannian metric on $\Sigma\subset\tilde{M}$ is non-negative and positive at some point. This Riemannian metric of $\Sigma$ is complete because it is pulled back from a compact manifold.
By the main theorem in [dCL] (see also [Sacksteder]), $\Sigma=\partial K\subset\mathbb{R}^{n}$ where $K\subset\mathbb{R}^{n}$ is convex closed subset in $\mathbb{R}^{n}$ with smooth boundary. Furthermore, $\Sigma$ is diffeomorphic to $S^{n-1}$ or to $\mathbb{R}^{n-1}$, hence it is the universal covering of a component of $\partial M$ with the induced metric. In the case $\Sigma\approx\mathbb{R}^{n-1}$, the Cheeger-Gromoll splitting theorem (Theorem 3 in [CG]) implies that the induced metric on $\Sigma$ is flat, which is a contradiction.
Hence $\Sigma\approx S^{n-1}$ and $K$ is a compact convex domain in $\mathbb{R}^{n}$ with smooth boundary. As $\tilde{M}$ is connected and complete, we have $\tilde{M}\subset K$ as a closed, hence compact subset. Assume $K\setminus\tilde{M}\neq\emptyset$. Then there exists a connected component $\Sigma^{\prime}\neq\Sigma$ of $\partial\tilde{M}$ and, again by [dCL], $\Sigma^{\prime}=\partial K^{\prime}$ for a convex closed subset $K^{\prime}\subset{\rm int}(K)\subset\mathbb{R}^{n}$. Since $\Sigma^{\prime}\subset\mathrm{int}(K)$, we have that $K^{\prime}\subset\mathrm{int}(K)$. In particular, $K^{\prime}\subset K$ and hence $K^{\prime}$ is compact. Arguing as before, we obtain $\tilde{M}\subset K^{\prime}$. Thus $\tilde{M}\subset\mathrm{int}(K)$, which is a contradiction.
Hence $\tilde{M}=K$ and $\tilde{M}$ is diffeomorphic to a ball. By the Brouwer fixed point theorem, we have $M=\tilde{M}$. This implies the assertion.
@@
Now assume $n=2$. Then the Gauss-Bonnet formula shows that the Euler characteristic of $M$ is positive, which implies that $M$ is diffeomorphic to a $2$-ball and, in particular, simply connected. Since the inclusion $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, we get $\tilde{M}^{\prime}=M^{\prime}$. In particular, $(M^{\prime},g^{\prime})$ is isometric to Euclidean space $\mathbb{R}^{2}$ and $(M,g)\subset(M^{\prime},g^{\prime})$ is isometric to a compact domain $K\subset\mathbb{R}^{2}$ with smooth, mean-convex boundary. This implies the assertion. ∎
@@ Appendix B An auxiliary estimate
Let $\Omega\subset\mathbb{R}^{n}$ be a compact, connected Lipschitz domain. Assume that $U$ is a non-empty open subset of $\Omega$. Suppose that $K>0$ is a given positive real number. Then
@@
This implies
$\int_{\Omega}(-F_{j})_{+}^{2}\to 0.$
Moreover, using the triangle inequality, we obtain
@@
After passing to a subsequence, the sequence $F_{j}$ converges to a function $F$ in $L^{2}(\Omega)$. The function $F$ satisfies $\int_{\Omega}F^{2}=1$. Moreover, $F=0$ almost everywhere in $U$, and $F\geq 0$ almost everywhere in $\Omega$.
Let $\eta:\mathbb{R}^{n}\to[0,\infty)$ denote a smooth function which is supported in the unit ball and satisfies $\int_{\mathbb{R}^{n}}\eta=1$. For each $\varepsilon>0$, we define a smooth function $\eta_{\varepsilon}:\mathbb{R}^{n}\to[0,\infty)$ by
$\eta_{\varepsilon}(x)=\frac{1}{\varepsilon^{n}}\eta\left(\frac{x}{\varepsilon}\right)$
@@
Let us fix an arbitrary point $p\in U$. Let $\varepsilon$ be a real number such that $0<\varepsilon<d(p,\partial U)$. Then $d(p,\partial\Omega)>\varepsilon$. We denote by $\Omega_{\varepsilon}$ the connected component of the set $\{x\in\Omega:d(x,\partial\Omega)>\varepsilon\}$ containing $p$. Let $F_{j}^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F_{j}$ with $\eta_{\varepsilon}$, and let $F^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F$ with $\eta_{\varepsilon}$. Then $|\nabla F_{j}^{\varepsilon}|\leq\eta_{\varepsilon}*|\nabla F_{j}|$ at each point in $\Omega_{\varepsilon}$. This implies
@@
at each point in $\Omega_{\varepsilon}$. We now pass to the limit as $j\to\infty$, keeping $\varepsilon$ fixed. Since $F_{j}\to F$ in $L^{2}(\Omega)$, we know that $F_{j}^{\varepsilon}\to F^{\varepsilon}$ and $|\nabla F_{j}^{\varepsilon}|\to|\nabla F^{\varepsilon}|$ at each point in $\Omega_{\varepsilon}$. Moreover, since $(|\nabla F_{j}|-KF_{j}\big{)}_{+}\to 0$ in $L^{2}(\Omega)$, it follows that $\eta_{\varepsilon}*(|\nabla F_{j}|-KF_{j})_{+}\to 0$ at each point in $\Omega_{\varepsilon}$. Putting these facts together, we conclude that
$|\nabla F^{\varepsilon}|-KF^{\varepsilon}\leq 0$
at each point in $\Omega_{\varepsilon}$. On the other hand, we know that $F=0$ almost everywhere in $U$. Since $d(p,\partial U)>\varepsilon$, it follows that $F^{\varepsilon}$ vanishes at the point $p$. Since $\Omega_{\varepsilon}$ is connected, standard ODE arguments imply that $F^{\varepsilon}=0$ at each point in $\Omega_{\varepsilon}$.

### v4

@@ Appendix A A metric characterization of the ball
  • The second fundamental form of $\partial M$ is weakly positive definite.
  • There exists a point in $\partial M$, where the second fundamental form is strictly positive definite.
@@
For $\varepsilon>0$ small enough, we consider the smooth manifold
@@
where $\sim$ identifies $\partial M\times(-\varepsilon,0]\subset\partial M\times(-\varepsilon,\infty)$ with an open neighborhood of $\partial M\subset M$ by the normal exponential map along $\partial M$ with respect to $g$. We denote the metric on $\partial M$ induced by $g$ by $g_{0}$ and the second fundamental form of $\partial M$ (considered as a $(1,1)$-tensor field) by $h$. The Gauss lemma, the Riccati equation for the second fundamental form and the vanishing of the curvature of $M$ imply that on $\partial M\times(-\varepsilon,0]$ the metric $g$ takes the form $g=dt\otimes dt+g_{t}$ where $g_{t}$ is the metric on $\partial M$ given by $g_{t}(X,Y)=g_{0}((\textup{id}+th)^{2}X,Y)$. Here $t\in(-\varepsilon,\infty)$ is the standard parameter. Since $h$ is weakly positive definite, we can use this formula to extend $g$ to $\partial M\times[0,\infty)$ and hence to a smooth metric $g^{\prime}$ on all of $M^{\prime}$. By [Baer-Gauduchon-Moroianu]*Theorem 7.2, the manifold $(M^{\prime},g^{\prime})$ is flat. It is easy to see that $(M^{\prime},g^{\prime})$ is complete.
[…]
We define $\tilde{M}=\pi^{-1}(M)$. Then $\tilde{M}$ is a closed domain in $\tilde{M}^{\prime}$ with smooth boundary. Since the embedding $M\hookrightarrow M^{\prime}$ is a homotopy equivalence, it follows that $\tilde{M}^{\prime}$ is connected. We distinguish two cases:
Case 1: Suppose that $n\geq 3$. By assumption, the second fundamental form of $\partial M$ is weakly positive definite at each point on $\partial M$, and is strictly positive definite at some point $p\in\partial M$. We denote by $\Sigma$ the connected component of $\partial M$ that contains the point $p$. Since $n\geq 3$, the Gauss equations imply that the sectional curvature of $\Sigma$ is nonnegative at each point on $\Sigma$, and is strictly positive at the point $p\in\Sigma$. Using the Cheeger-Gromoll splitting theorem [Cheeger-Gromoll]*Theorem 3, we conclude that every covering of $\Sigma$ is compact.
In the next step, we fix a point $\tilde{p}\in\partial\tilde{M}$ with $\pi(\tilde{p})=p$. We denote by $\tilde{\Sigma}$ the connected component of $\partial\tilde{M}$ that contains the point $\tilde{p}$. Then $\tilde{\Sigma}$ is a covering of $\Sigma$. In view of the discussion above, $\tilde{\Sigma}$ is compact.
[…]
If $M$ is not simply connected, then there exists an isometry of $\tilde{M}^{\prime}$ which maps $\tilde{M}$ to itself and has no fixed points. This contradicts the Brouwer fixed point theorem. Thus, $M$ is simply connected, and so is $M^{\prime}$. From this, the assertion follows.
@@
Case 2: Suppose that $n=2$. In this case, the Euler characteristic of $M$ is strictly positive by the Gauss-Bonnet theorem. Since $M$ is connected, it follows that $M$ is diffeomorphic to a disk $B^{2}$. In particular, $M$ is simply connected, and so is $M^{\prime}$. Consequently, $(M^{\prime},g^{\prime})$ is isometric to $\mathbb{R}^{2}$ and $(M,g)$ can be identified with a compact convex domain in $R^{2}$ with smooth boundary. ∎
@@ Appendix B An auxiliary estimate
Let $\Omega\subset\mathbb{R}^{n}$ be a compact connected Lipschitz domain. Assume that $U$ is a nonempty open subset of $\Omega$. Suppose that $K>0$ is a given positive real number. Then
@@
Using the triangle inequality, we obtain
@@
After passing to a subsequence, the sequence $F_{j}$ converges in $L^{2}(\Omega)$ to a nonnegative function $F$. The function $F$ satisfies $\int_{\Omega}F^{2}=1$. Moreover, $F=0$ almost everywhere in $U$.
Let $\eta\colon\mathbb{R}^{n}\to[0,\infty)$ denote a smooth function which is supported in the unit ball and satisfies $\int_{\mathbb{R}^{n}}\eta=1$. For each $\varepsilon>0$, we define a smooth function $\eta_{\varepsilon}\colon\mathbb{R}^{n}\to[0,\infty)$ by
$\eta_{\varepsilon}(x)=\varepsilon^{-n}\,\eta(\varepsilon^{-1}x)$
@@
Let us fix an arbitrary point $p\in U$. Let $\varepsilon$ be a real number such that $0<\varepsilon<d(p,\partial U)$. Then $d(p,\partial\Omega)>\varepsilon$. We denote by $\Omega_{\varepsilon}$ the connected component of the set
$\{x\in\Omega\colon d(x,\partial\Omega)>\varepsilon\}$
containing $p$. Let $F_{j}^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F_{j}$ with $\eta_{\varepsilon}$, and let $F^{\varepsilon}\colon\Omega_{\varepsilon}\to\mathbb{R}$ denote the convolution of $F$ with $\eta_{\varepsilon}$. Then $|\nabla F_{j}^{\varepsilon}|\leq\eta_{\varepsilon}*|\nabla F_{j}|$ at each point in $\Omega_{\varepsilon}$. This implies
@@
at each point in $\Omega_{\varepsilon}$. We now pass to the limit as $j\to\infty$, keeping $\varepsilon$ fixed. Since $F_{j}\to F$ in $L^{2}(\Omega)$, we know that $F_{j}^{\varepsilon}\to F^{\varepsilon}$ and $|\nabla F_{j}^{\varepsilon}|\to|\nabla F^{\varepsilon}|$ at each point in $\Omega_{\varepsilon}$. Moreover, since $(|\nabla F_{j}|-KF_{j}\big)_{+}\to 0$ in $L^{2}(\Omega)$, it follows that $\eta_{\varepsilon}*(|\nabla F_{j}|-KF_{j})_{+}\to 0$ at each point in $\Omega_{\varepsilon}$. Putting these facts together, we conclude that $|\nabla F^{\varepsilon}|-KF^{\varepsilon}\leq 0$ at each point in $\Omega_{\varepsilon}$. On the other hand, we know that $F=0$ almost everywhere in $U$. Since $d(p,\partial U)>\varepsilon$, it follows that $F^{\varepsilon}$ vanishes at the point $p$. Since $\Omega_{\varepsilon}$ is connected, standard ODE arguments imply that $F^{\varepsilon}=0$ at each point in $\Omega_{\varepsilon}$.